Crew 200 Sol Summaries

From Marspedia
Jump to navigation Jump to search


POSTED ON DECEMBER 15, 2018 Sol Summary – December 15th Crew 200 Sol Summary Report 15-DEC-2018

Sol: 13

Summary Title: Cleaning, gathering food, and two happy crews meet

Author’s name: Crew 200 and 201

Mission Status: Transitioning the habitat from Crew 200 to Crew 201.

Sol Activity Summary: One crew 200 members drove to Grand Junction this morning to pick up Crew 201! Two members from crew 200 remained in the habitat and thoroughly cleaned. Crew 200 welcomed and oriented Crew 201 and also helped them gather new food from Hanksville. A lot of delicious, dehydrated food! Both crews will enjoy dinner together tonight!

Look Ahead Plan: Crew 200 will leave for Grand Junction tomorrow and Crew 201 will eagerly begin their mission!

Anomalies in work: N/A

Weather: Chilly, snow still melting, muddy

Crew Physical Status: All crew are happy and healthy

EVA: N/A

Reports to be filed: Operations Report, GreenHab Report, Sol Report,

Support Requested: N/A

POSTED ON DECEMBER 14, 2018 Sol Summary – December 14 Crew 200 Sol Summary Report 14-DEC-2018

Sol: 12

Summary Title: Driving to the moons and cleaning the habitat

Author’s name: Lindsay Rutter

Mission Status: Simulation ended at 1:00PM today. All crew members happy and healthy.

Sol Activity Summary: We spent much of the morning finishing our reports about this week. Each day, we have been documenting our observations of the dynamics of a 3 member crew. Today, we summarized it over a decadent dessert. One crew member made cinnamon chocolate cake by modifying “Kyle’s Carrot Cake”, a recipe in the habitat cookbook “Mars Home Cooking: Recipes from the Mars Desert Research Station”. After we discontinued our simulation as planned, we drove two rovers to an array of “moons” north of the habitat. White Moon. Beige Moon. Yellow Moon. Gray Moon.

Look Ahead Plan: Tonight and tomorrow, we will be thoroughly cleaning the habitat and its surrounding buildings. Tomorrow, we will also drive to Hanksville to pick up the next crew and drive them to MDRS. We will introduce the MDRS to them and have discussions with them about the rules and expectations of conducting missions here.

Anomalies in work: None reported

Weather: Snow melting; chilly with moderate cloud coverage.

Crew Physical Status: No crew members reported physical problems today.

EVA: N/A

Reports to be filed: Operations Report, GreenHab Report, Pictures Report, Journalist Report, Sol Report, Mission Summary Report

Support Requested: N/A

POSTED ON DECEMBER 13, 2018 Sol Summary – December 13th Sol: 11

Summary Title: The first 3 crew EVA on Mars

Author’s name: Makiah Eustice

Mission Status: Crew conducted an EVA to the North of Stream Bed connector.

Sol Activity Summary: Crew worked on projects in the morning, waiting to go on EVA after a visit

from guests. We enjoyed home-made bread from Antoine throughout the day. We made the decision to

end out sim tomorrow, which mean we are wrapping up our projects and making grand last meals.

Look Ahead Plan: Tomorrow we will wrap up our projects and end simulation after lunch.

Anomalies in work: None

Weather: Cold, but drier overall.

Crew Physical Status: No crew members reported physical problems today.

EVA: EVA Crew went by Rover to Galileo road, and then just past Reservoir Dam to investigate

comms capability.

Reports to be file: Operations Report, EVA Report, GreenHab Report, Pictures Report, Journalist

Report, Science Report

Support Requested: N/A

POSTED ON DECEMBER 12, 2018 Sol Summary – December 12th Crew 200 Sol Summary 12Dec2018

Crew 200 Sol Summary Report 12-DEC-2018

Sol: 10

Summary Title: Exploring Phobos Peak and examining unusual objects collected during EVA

Author’s name: Lindsay Rutter

Mission Status: All crew members are happy and healthy

Sol Activity Summary: Each morning, the first object I see when I look out the east window is the silhouette of Phobos Peak Rock as the rising sun lights it up from behind. We had a beautiful sunrise this morning and Phobos Peak Rock emerged all the more majestic than usual. I was happy to remember that it was our approved geographical destination for our EVA. We conducted another 3 crew EVA and walked to Phobos Peak. We then explored its surrounding areas, discovering a curious aluminum object near the peak, which we returned to the habitat for investigation. Winds became violent shortly after our return, but we celebrated our successful EVA safely inside our habitat with a dish of quinoa peppered with oregano from the GreenHab.

Look Ahead Plan: We will continue conducting 3 crew EVAs and mapping radio strength and connectivity. We will also continue our collaborative projects.

Anomalies in work: Heavy winds caused minor damage in the tunnels, which we repaired. Heating system in the habitat has been working on and off; it appears to be working now but we will continue monitoring it.

Weather: Snow melting; clear in the morning but clouds and heavy winds developed in the early afternoon.

Crew Physical Status: No crew members reported physical problems today.

EVA: A walking EVA was conducted at and around Phobos Peak. We continued our communication protocol to assess radio strength and connectivity in areas around the habitat. Our EVA was almost exactly two hours.

Reports to be file: Operations Report, EVA Report, EVA Request, GreenHab Report, Pictures Report, Journalist Report, Sol Report

Support Requested: N/A

POSTED ON DECEMBER 11, 2018 Sol Summary – December 11th Sol: 9

Summary Title: First 3 crew EVA using rovers.

Author’s name: Lindsay Rutter

Mission Status: All crew members are happy after 3 crew EVA using rovers.

Sol Activity Summary: We conducted a 3 crew EVA today using rovers and continued mapping communication strength and readibility. We realized yesterday that we did not document any photos during our first 3 crew EVA because we were concentrating on our communication protocol. Today, we reserved extra time at the beginning and end of our EVA to better document it with photos. After our EVA, we added our newest observations about 3 crew EVAs to a list we are keeping this week and will share at the end of our mission. We continued plotting the radio test values we collected during our EVA onto maps and collaborating with projects in the habitat. We also caught a habitat mouse today; tomorrow, we will remove it a significant distance from the habitat as per protocol. We will be careful to monitor and report any other problematic pests around the habitat.

Look Ahead Plan: We plan to continue mapping the radio connectivity between EVA crew and HabCom in northern and southern areas and testing how to safely and effectively conduct 3 crew EVAs. We will also continue plotting our tested radio connectivity values onto maps together and collaborating between projects in the habitat and GreenHab.

Anomalies in work: Mouse was found in a trap inside the habitat. We will follow protocol and remove it during our EVA tomorrow.

Weather: Cold, muddy from snow melt.

Crew Physical Status: No crew members reported physical problems today.

EVA: Crew completed an EVA with 2 members outside and 1 member in the habitat. EVA members performed an engineering check, drove north with Curiosity to Cowboys’ corner (this was further than intended due to recent snow melting obstructing landmarks), returned to the habitat, switched rovers, drove south with Opportunity to Robert’s Rock Garden, walked along Kissing Camel Ridge, returned to the habitat, took several photos to document our 3 crew EVA. We applied our communication and GPS coordinate relay system throughout the EVA.

Reports to be filed: Operations Report, EVA Report, EVA Request, GreenHab Report, Pictures Report, Journalist Report, Sol Report

Support Requested: N/A

POSTED ON DECEMBER 10, 2018 Sol Summary – December 10th Sol:8

Summary Title: The first 3 crew EVA on Mars

Author’s name: Makiah Eustice

Mission Status: Crew conducted an EVA with 2 crew members out and only one member on base, mopped the Science Dome.

Sol Activity Summary: Crew conducted first 3-crew member EVA to explore and map the regions of communication strength. We planned contingency plans for loss of signal and navigation. Afterwards, the crew mapped the locations of perceived path by each crew member against the estimated path based on recorded GPS coordinates. The rest of the day was relaxed and quiet, taking time to clean the mudded floor of the Science Dome and have computer-based work on our projects.

Look Ahead Plan: Tomorrow an driving/walking EVA is planned. Crew projects will continue collaborative projects.

Anomalies in work: The observatory was open most of the day. It closed once we restarted the generator.

Weather: Cold, muddy from snow melt.

Crew Physical Status: No crew members reported physical problems today.

EVA: EVA crew went, by foot, to Kissing Camel Ridge East to explore the limits of walkie talkie comms in certain regions. Comms checks were performed every minute between HabCom and EVA, and GPS coordinate was exhanged every few minutes. EVA was planned for 2 hours and was conducted in exactly 2 hours.

Reports to be file: Operations Report, EVA Report, EVA Request, GreenHab Report, Pictures Report, Journalist Report, HSO Report

Support Requested: N/A

POSTED ON DECEMBER 9, 2018 Sol Summary – December 9th Crew 200 Sol Summary Report 9-Dec-2018

Sol: 7

Summary Title: Reduced crew EVA planning, replenishing supplies, and fixing tunnels

Author’s name: Lindsay Rutter

Mission Status: Nominal

Sol Activity Summary: This morning we rejuvenated a bit. We ate breakfast two hours past our usual meal time. We had a long and engaging conversation with Shannon on a variety of topics, including how to conduct EVAs during this reduced-crew mission. As today was our transition day between our nominal and extended missions, we returned to Hanskville to replenish gas and water and to throw out the non-burnable trash. After that, we fixed some of the tunnels connecting to the RAM and GreenHab to make sure they are extra safe. We can’t wait to to return to Mars this evening for our extended mission!

Look Ahead Plan: We plan to safely conduct our first reduced-crew EVA tomorrow and begin to map out the radio communication zones around the habitat. We will also keep track of what methods work best for thoroughly cleaning the ScienceDome floor. We brainstormed numerous project ideas today and will begin to focus on the most practical and meaningful ones starting tomorrow.

Anomalies in work: No anomalies.

Weather: Clear, sunny, and cool

Crew Physical Status: Nominal

EVA: N/A

Reports to be filed: N/A

Support Requested: None

DECEMBER 8, 2018 Sol Summary – December 8th Crew 200 Sol Summary Report 08-DEC-2018

Sol: 6

Summary Title: First solar observation, mission extended for reduced crew members

Author’s name: Lindsay Rutter

Mission Status: Mission is completed. A proposal was made to extend the mission with a reduced crew number of three people who will work toward a relatively new set of objectives. Proposal was approved.

Sol Activity Summary: We woke up today to the smell of lobster biscuits and herb garlic butter. Several crew members had been salivating over that box of lobster biscuits since Sol 1, so it was a celebratory meal that marked the end of our nominal mission. The skies were clear today and our crew astronomer was kind enough to show several crew members how to use the solar observatory. After cleaning and bag packing, we drove to Hanksville to share lunch together and part ways. Four crew members continued to Grand Junction, while three crew members returned to the habitat. It was hard to see the crew split into half, but the three remaining crew members are determined to use what they have learned from the full team to make the extended mission worthwhile at both a personal level and team level.

Look Ahead Plan: The reduced crew members will plan new and modified research projects and brainstorm possible reduced crew EVAs.

Anomalies in work: N/A

Weather: Warmer than recently, snow melting, clear and sunny skies.

Crew Physical Status: No crew members reported physical problems today.

EVA: N/A

Reports to be file: N/A

Support Requested: N/A

POSTED ON DECEMBER 7, 2018 Sol Summary – December 7th Sol: 5

Summary Events: Final day of Sim for Whole Crew. Day focused on completing projects

Author’s name: Oakley Jennings-Fast

Mission Status: Nominal

Sol Activity Summary: Engineering check completed with Crew Engineer and Crew Astronomer. Each crew member focused on completing their projects. The crew members, who completed their projects, aided the others in their projects. Crew Astronomer taught two of the crew members who are staying for a second week of sim how to run the Solar Observatory. Fresh lettuce was harvested for the crews’ final meal all together on sim.

Look Ahead Plan: Midnight tonight, we will end sim. Tomorrow midday, 4 crew members will go back to Grand Junction to fly home on Sunday. Three crew members are staying an additional week and have coordinated with Shannon.

Anomalies in work: None.

Weather: Cloudy, no precipitation

Crew Physical Status: Nominal

EVA: None

Reports to be filed: N/A

Support Requested: None

POSTED ON DECEMBER 6, 2018 Mission Support Crew 200 Sol Summary December 6th Sol: 4

Summary Title: A productive day of science, EVAs, visitors, and outreach.

Author’s name: Lindsay Rutter

Mission Status: All crew members happy, healthy, and productive.

Sol Activity Summary: Crew ate all meals together. Science projects continued throughout the day. Two EVAs were conducted. Crew members welcomed more visitors to the habitat. Group photographs were taken with crew members holding banners from girl scout troupe from Qatar.

Look Ahead Plan: No EVAs tomorrow; crew members will focus on research projects.

Anomalies in work: None reported.

Weather: Cold with snow and mud on the ground.

Crew Physical Status: No crew members reported physical problems today.

EVA: Early EVA consisted of four female crew members and was led by female crew commander. Destination was about one mile north of the habitat. Soil samples were collected. Later EVA consisted of three crew members with media guests. Destination was hill next to the habitat.

Reports to be file: N/A

Support Requested: N/A

POSTED ON DECEMBER 4, 2018 Sol Summary – December 4th Sol: 2

Summary Title: Safety training, middle school outreach, and science experiments

Author’s name: Lindsay Rutter

Mission Status: Crew practiced emergency training, science experiments underway, public outreach underway

Sol Activity Summary: We discussed and acted out various emergency/health scenarios, such as strokes, electric shock, hypothermia, fire, and CPR cases. Safety checklist was completed in regards to fire alarms, CO alarms, and medical kits. Microscopes were inspected. Questions about living on MDRS and Mars from eighth graders were answered. Personal experiments were underway.

Look Ahead Plan: Weather-pending EVAs for drone mapping and infrared camera shots will be conducted. Science experiments will continue.

Anomalies in work: No astronomy possible due to cloudy conditions.

Weather: Cold with snow and mud on the ground.

Crew Physical Status: No crew members reported physical problems today.

EVA: N/A

Reports to be file: N/A

Support Requested: N/A

POSTED ON DECEMBER 3, 2018 Sol Summary – December 3rd Health and Safety Officer Report Lindsay Rutter, Health and Safety Officer of Crew 200

Crew 200 Sol Summary Report 03Dec2018

Sol: 1 Summary Title: Introduced guests and fixed solar observatory Author’s name: Lindsay Rutter Mission Status: All crew members are healthy and productive. Sol Activity Summary: We interacted with guests and fixed solar observatory. Look Ahead Plan: We will do science most of the day tomorrow. Anomalies in work: Solar observatory had issues that needed fixing. Weather: Unexpected snowy weather created muddy and slippery ground. Crew Physical Status: One crew member felt headache due to cold weather, rested, and has no more symptoms. EVA: No EVAs today due to inclement weather. Reports to be file: Support Requested:


Source